OpenStudy (anonymous):

What is the simplified form of each expression? 5(14 - 2)2 ÷ 2

OpenStudy (callisto):

\[5(14 - 2)^2 ÷ 2 = 5(12)^2 / 2 = 5(144)/2 = 5(72) = ?\]

OpenStudy (anonymous):

ummmm........ each expression?? the only thing we can get from what you've given to us is 60??? is the 2 after the brackets an indice??

OpenStudy (callisto):

I assumed that was a power 2 :|

OpenStudy (anonymous):

yeah i got 60 to but whats the question was asking was to solve the equation and i was thinking i had the answer wrong but the 2 behind the bracket was squared thanks guys!!!

OpenStudy (callisto):

equation?! there should be an unknown?!

OpenStudy (anonymous):

^WORD.

OpenStudy (anonymous):

5(12) x 2÷2 = (60 x 2) ÷ 2 = 120 ÷ 2 =60. Think this is right.

OpenStudy (anonymous):

What's an 'indice'?

OpenStudy (anonymous):

a power. like x^2